OpenStudy (brinazarski):

Can someone help me out with this Trig question? When sinx = -8/17 and x lies in Quadrant III and cosy = -4/5 and y lies in Quadrant II, what is cos(x-y?) So far, I did cos(-8/17 - -4/5) = cos-8/17cos-4/5 + sin-8/17sin-4/5 And now I'm stuck.

OpenStudy (anonymous):

we need \[\cos(x-y)=\cos(x)\cos(y)+\sin(x)\sin(y)\] so we want \[\cos(x), \sin(x), \cos(y), \sin(y)\]

OpenStudy (brinazarski):

I think so.

OpenStudy (anonymous):

\[\sin(x)=-\frac{8}{17}\]

OpenStudy (anonymous):

sorry that was a type, but before we start let me point out something i think is confusing youi

OpenStudy (anonymous):

the sine of x IS -8/17, you do not want the sine of -8/17 you just need these numbers, not the sine or cosine of these numbers

OpenStudy (brinazarski):

cos(x) = -15/17, you wanted that earlier, right?

OpenStudy (anonymous):



OpenStudy (brinazarski):

So... I'm not supposed to be using the formula I used...?

OpenStudy (anonymous):

yes, but you are confused i think about the formula. you need the numbers themselves, not the cosine and sine of the numbers

OpenStudy (anonymous):

in other words just work with the numbers \[\sin(x)=-\frac{8}{15}\] \[\cos(x)=-\frac{15}{17}\] \[\cos(y)=-\frac{4}{2}\] \[\sin(y)=\frac{3}{5}\]

OpenStudy (anonymous):

typo there \[\cos(y)=-\frac{4}{5}\] of course now plug the numbers directly into the formula

OpenStudy (anonymous):

and in quadrant II sine is positive

OpenStudy (anonymous):

so your answer is \[-\frac{8}{17}\times -\frac{4}{5}+(-\frac{15}{17})\times \frac{4}{5}\]

OpenStudy (anonymous):

clear right? those are the sines and cosines, you do not take the sines and cosines of those numbers

OpenStudy (brinazarski):

Uh... I'm a little confused, hold on a sec x_x;;;

OpenStudy (anonymous):

ok i will wait. this is a common mistake, i have seen it before

OpenStudy (anonymous):

point is that \[\cos(x)=-\frac{3}{5}\] you do not take the cosine of that number. that number is your cosine

OpenStudy (brinazarski):

If it's cos(x)cos(y), shouldn't it be -15/17 x -4/5?

OpenStudy (anonymous):

yes sorry that was a typo on my part

OpenStudy (anonymous):

you are right just don't do this cos-8/17cos-4/5 + sin-8/17sin-4/5 instead do this -8/17 *-4/5 + -8/17*4/5

OpenStudy (brinazarski):

According to my calculator, that's zero, unless there's another typo ><....

OpenStudy (anonymous):

no it is not zero